A dry cell of emf 1.6V and internal resistance 0.1ohm is connected to a resistor of R Ohms. if 2A current is flowing in the circuit, calculate the voltage drop across R

Grade:12

1 Answers

utpal raj
7 years ago
hiii anusha 
total resistance in the circuit = R+0.1=R1
according to ohms law V=IR1
1.6=2*R1
R1=0.8 ohm
0.8 = 0.1+R 
R=0.7 ohm
voltage across R = 0.7*2=1.4 volt , 
the current passing through R is 2 amp 
so , hope this clears 
please aproove 
good luck